A hydraulic press has a small piston of diameter 2 cm. If a force of 60 N is applied to the smaller piston, a force of 300 N is exerted on the larger piston. What is the diameter of the larger piston?


1
Expert's answer
2021-05-07T08:11:47-0400

Pascal's Principle: "F1\/A1=F2\/A2"


F1=60N r1=(2÷2)1/100


r1=0.01m


F2=300N r2=?


A=(πr2)


A1=π(0.012)m2


A2=π(r2)m2


(60N/π(0.01)2)=300N/π(r2)


r2/0.012=5


√ r2=√0.0005


r=0.0224m


r= (0.224x100) cm


r=2.24cm


Diameter= (2.24cm x2)


D=4.48 cm